### v3

# An Interpolation of Metrics and Spaces of Metrics

Abstract.
As a generalization of Hausdorff’s extension theorem of metrics, we prove an interpolation theorem of a family of metrics defined on closed subsets of metrizable spaces. As an application, we investigate typicality of subsets of moduli spaces of metrics. We observe that various sets of all metrics with properties appearing in metric geometry are dense intersections of countable open subsets in spaces of metrics on metrizable spaces. For instance, our study is applicable to the set of all non-doubling metrics and the set of all non-uniformly disconnected metrics.
Key words and phrases:
Metric space, Baire space. Space of metrics
2020 Mathematics Subject Classification:
Primary 54E35; Secondary 54E52
1. Introduction
1.1. Backgrounds
In 1930, Felix Hausdorff [14] proved the extension theorem stating that for every metrizable space $X$, for every closed subset $A$ of $X$ and for every metric $d$ on $A$ generating the same topology of $A$, there exists a metric $D$ on $X$ such that $D$ generates the same topology of $X$ and $D|_{A^{2}}=d$ (see Theorem 2.1). Motivated by this result, the author thought that Hausdorff’s theorem suggested the study of spaces of metrics and tried to investigate their topological properties.
For a metrizable space $X$, we denote by $\mathrm{Met}(X)$ the set of all metrics on $X$ generating the same topology of $X$. We define a function $\mathcal{D}_{X}:\mathrm{Met}(X)\times\mathrm{Met}(X)\to[0,\infty]$ by
$\mathcal{D}_{X}(d,e)=\sup_{(x,y)\in X^{2}}|d(x,y)-e(x,y)|.$
The function $\mathcal{D}_{X}$ is a metric on $\mathrm{Met}(X)$ valued in $0,\infty]$. Throughout this paper, we consider that $\mathrm{Met}(X)$ is always equipped with the topology generated by $\mathcal{D}_{X}$. To investigate topological properties of $\mathrm{Met}(X)$, we first generalize the Hausdorff extension theorem to an interpolation theorem of metrics with an approximation by $\mathcal{D}_{X}$ (Theorem 1.1 or 3.4). As applications of it, for a certain property $\mathcal{P}$ on metric spaces, we prove that the set set $\{\,d\in\mathrm{Met}(X)\mid\text{$(X,d)$ satisfies $\mathcal{P}$}\,\}$ is generic in $\mathrm{Met}(X)$ (see Theorems 1.2 and 1.3). This result can be considered as an analogue of Banach’s famous result, which states that in the continuous function space on $[0,1]$, a generic function is nowhere differentiable.
Since a metric on $X$ determines the geometry on $X$, studying the moduli space $\mathrm{Met}(X)$ of metrics is equivalent to investigating the moduli of geometries that can be expanded on the space $X$.
1.2. Main results
1.2.1. Interpolation of metrics
A family $\{S_{i}\}_{i\in I}$ of subsets of a topological space $X$ is said to be discrete if for every $x\in X$ there exists a neighborhood of $x$ intersecting at most single member of $\{S_{i}\}_{i\in I}$.
Our first main results is the following interpolation theorem:
Theorem 1.1.
Let $X$ be a metrizable space, and let $\{A_{i}\}_{i\in I}$ be a discrete family of closed subsets of $X$. Then for every metric $d\in\mathrm{Met}(X)$, and for every family $\{e_{i}\}_{i\in I}$ of metrics with $e_{i}\in\mathrm{Met}(A_{i})$, there exists a metric $m\in\mathrm{Met}(X)$ satisfying the following:
  1. (1)
    for every $i\in I$ we have $m|_{A_{i}^{2}}=e_{i}$;
  2. (2)
    $\mathcal{D}_{X}(m,d)=\sup_{i\in I}\mathcal{D}_{A_{i}}(e_{A_{i}},d|_{A_{i}^{2}})$.
Moreover, if $X$ is completely metrizable, and if each $e_{i}\in\mathrm{Met}(A_{i})$ is a complete metric, then we can choose $m\in\mathrm{Met}(X)$ as a complete one.
Remark 1.1.
For every metrizable space $X$, and for every closed subset $A$ of $X$, Nguyen Van Khue and Nguyen To Nhu [20] constructed a Lipschitz metric extensor from $(\mathrm{Met}(A),\mathcal{D}_{A})$ into $(\mathrm{Met}(X),\mathcal{D}_{X})$, and a monotone continuous metric extensor from $\mathrm{Met}(A)$ into $\mathrm{Met}(X)$; moreover, if $X$ is completely metrizable, then each of these metric extensors maps any complete metric in $\mathrm{Met}(A)$ into a complete metric in $\mathrm{Met}(X)$. To obtain such metric extensors, they used the Dugundji extension theorem concerning locally convex topological linear spaces. For more information of simultaneous extensions of metrics, see [3], [5], [4], [29], [25], [26], [24], and [17]. It seems difficult to prove Theorem 1.1 using these extension result because they do not preserve a given metric $d$.
A central idea of the proof of Theorem 1.1 is a correspondence between a metric on a metrizable space and a topological embedding from a metrizable space into a Banach space. A metric $d$ on a metrizable space $X$ induces a topological embedding from $X$ into a Banach space such as the Kuratowski embedding (see Theorem 2.6). Conversely, a topological embedding $F$ from a metrizable space $X$ into a Banach space $V$ with norm $\|\cdot\|_{V}$ induces a metric $m\in\mathrm{Met}(X)$ on $X$ defined by $m(x,y)=\|F(x)-F(y)\|_{V}$. In the proof of Theorem 1.1, we utilize this correspondence to translate the statement of Theorem 1.1 into an approximation problem on topological embeddings into a Banach space. We then resolve such a problem by using the Michael continuous selection theorem (see Theorem 2.2), and by using a similar method to Kuratowski [21] (see also [15] and [1]) of converting a continuous function into a topological embedding by extending a codomain.
1.2.2. Moduli spaces of metrics
Theorem 1.1 enables us to investigate dense $G_{\delta}$ subsets in the topology of the space $(\mathrm{Met}(X),\mathcal{D}_{X})$ for a metrizable space $X$. To describe our second result precisely, we define a class of geometric properties that unifies various properties appearing in metric geometry. Due to this purpose, the definition may seem somewhat complicated at first glance.
Let $\mathcal{P}^{*}(\mathbb{N})$ be the set of all non-empty subsets of $\mathbb{N}$. For a topological space $T$, we denote by $\mathcal{F}(T)$ the set of all closed subsets of $T$. For a subset $W\in\mathcal{P}^{*}(\mathbb{N})$, and for a set $S$, we denote by $\operatorname{Seq}(W,S)$ the set of all finite injective sequences $\{a_{i}\}_{i=1}^{n}$ in $S$ with $n\in W$.
Definition 1.1.
Let $Q$ be an at most countable set, $P$ a topological space. Let $F:Q\to\mathcal{F}(P)$ and $G:Q\to\mathcal{P}^{*}(\mathbb{N})$ be maps. Let $Z$ be a set. Let $\phi$ be a correspondence assigning a pair $(q,X)$ of $q\in Q$ and a metrizable space $X$ to a map $\phi^{q,X}:\operatorname{Seq}(G(q),X)\times Z\times\mathrm{Met}(X)\to P$. We say that a sextuple $(Q,P,F,G,Z,\phi)$ is a transmissible paremeter if for every metrizable space $X$, for every $q\in Q$, and for every $z\in Z$ the following are satisfied:
  1. (TP1)
    for every $a\in\operatorname{Seq}(G(q),X)$, and for every $z\in Z$, the map $\phi^{q,X}(a,z):\mathrm{Met}(X)\to P$ defined by $\phi^{q,X}(a,z)(d)=\phi^{q,X}(a,z,d)$ is continuous;
  2. (TP2)
    for every $d\in\mathrm{Met}(X)$, if $S$ is a subset of $X$ and $a\in\operatorname{Seq}(G(q),S)$, then we have $\phi^{q,X}(a,z,d)=\phi^{q,S}(a,z,d|_{S^{2}})$.
The diameter $\phi^{q,X}(a,z,X)=\delta_{d}(\{a_{i}\}_{i=1}^{n})$ is a typical example of a map appearing in the condition (TP1).
We introduce a property determined by a transmissible parameter.
Definition 1.2.
Let $\mathfrak{G}=(Q,P,F,G,Z,\phi)$ be a transmissible parameter. We say that a metric space $(X,d)$ satisfies the $\mathfrak{G}$-transmissible property if there exists $q\in Q$ such that for every $a\in\operatorname{Seq}(G(q),X)$ and for every $z\in Z$ we have $\phi^{q,X}(a,z,d)\in F(q)$. We say that $(X,d)$ satisfies the anti-$\mathfrak{G}$-transmissible property if $(X,d)$ satisfies the negation of the $\mathfrak{G}$-transmissible property; namely, for every $q\in Q$ there exist $a\in\operatorname{Seq}(G(q),X)$ and $z\in Z$ with $\phi^{q,X}(a,z,d)\in X\setminus F(q)$. A property on metric spaces is a transmissible property (resp. anti-transmissible property) if it is equivalent to a $\mathfrak{G}$-transmissible property (resp. anti-$\mathfrak{G}$-transmissible property) for some transmissible parameter $\mathfrak{G}$.
The class of transmissible properties contains various properties appeared in metric geometry.
Example 1.1.
The following properties on metric spaces are transmissible properties (see Section 4).
  1. (1)
    the doubling property;
  2. (2)
    the uniform disconnectedness;
  3. (3)
    satisfying the ultratriangle inequality;
  4. (4)
    satisfying the Ptolemy inequality;
  5. (5)
    the Gromov $\operatorname{Cycl}_{m}(0)$ condition;
  6. (6)
    the Gromov hyperbolicity.
To state our second result, we need a more additional condition for transmissible properties.
Definition 1.3.
Let $\mathfrak{G}=(Q,P,F,G,Z,\phi)$ be a transmissible parameter. We say that $\mathfrak{G}$ is singular if for each $q\in Q$ and for every $\epsilon\in(0,\infty)$ there exist $z\in Z$ and a finite metrizable space $(L,d_{L})$ such that
  1. (1)
    $\delta_{d_{L}}(L)\leq\epsilon$, where $\delta_{d_{L}}(L)$ stands for the diameter of $L$;
  2. (2)
    $\operatorname{card}(L)\in G(q)$, where $\operatorname{card}$ stands for the cardinality;
  3. (3)
    $\phi^{q,L}(L,z,d_{L})\in X\setminus F(q)$.
As we will see in Lemma 4.5, the singularity of a transmissible parameter $\mathfrak{G}$ is equivalent to the existence of a special countable metric subspace that satisfies the anti-$\mathfrak{G}$-transmissible property.
Note that not all transmissible parameters are singular; especially, the Gromov hyperbolicity does not have a singular transmissible parameter (see Proposition 4.22).
Due to Theorem 1.1, we obtain the second main result on dense $G_{\delta}$ subsets in spaces of metrics:
Theorem 1.2.
Let $\mathfrak{G}$ be a singular transmissible parameter. For every non-discrete metrizable space $X$, the set of all $d\in\mathrm{Met}(X)$ for which $(X,d)$ satisfies the anti-$\mathfrak{G}$-transmissible property is dense $G_{\delta}$ in $\mathrm{Met}(X)$.
Remark 1.2.
Theorem 1.2 holds true for the space $\operatorname{Comp}(X)$ of all complete metrics in $\mathrm{Met}(X)$ (see Theorems 4.7).
We can apply Theorem 1.2 to the properties (1)(5) mentioned in Example 1.1. Therefore we conclude that the set of all metrics not satisfying these properties are dense $G_{\delta}$ in spaces of metrics (see Theorem 4.12 and Corollary 4.17). We also conclude that the set of all metrics with rich pseudo-cones is dense $G_{\delta}$ in spaces of metrics (see Theorem 4.15).
Our third result is based on the fact that for every second-countable locally compact space $X$, the space $\mathrm{Met}(X)$ is Baire (see Lemma 5.1). For a property $P$ on metric spaces, we say that a metric space $(X,d)$ satisfies the local $P$ if every non-empty open metric subspace of $X$ satisfies the property $P$.
As a local version of Theorem 1.2, we obtain the following:
Theorem 1.3.
Let $X$ be a second-countable, locally compact locally non-discrete space. Then for every singular transmissible parameter $\mathfrak{G}$, the set of all metrics $d\in\mathrm{Met}(X)$ for which $(X,d)$ satisfies the local anti-$\mathfrak{G}$-transmissible property is dense $G_{\delta}$ in $\mathrm{Met}(X)$.
Recall that all second-countable locally compact spaces are metrizable, which is a consequence of the Urysohn metrization theorem.
1.2.3. Organization
This paper is organized as follows. In Section 2, we review the basic and classical theorems on topological spaces and metric spaces. Section 3 is devoted to the proof of Theorem 1.1. In Section 4, we prove Theorem 1.2 and show that various properties in metric geometry are transmissible properties. Section 5 contains the proof that $\mathrm{Met}(X)$ is Baire when $X$ is second-countable and locally compact. We also show Theorem 1.3.
Acknowledgements.
The author would like to thank Professor Koichi Nagano for his advice and constant encouragement. This work was supported by JSPS KAKENHI Grant Number 18J21300, and partially supported by JSPS KAKENHI Grant Number JP24KJ0182. The author also would like to thank the referee for variable and helpful comments to improve the present paper.
2. Preliminaries
In this paper, the symbol $\mathbb{N}$ stands for the set of all positive integers.
Let $(X,d)$ be a metric space. Let $A$ be a subset of $X$. We denote by $\delta_{d}(A)$ the diameter of $A$. We denote by $B(x,r)$ (resp. $U(x,r)$) the closed (resp. open) ball centered at $x$ with radius $r$.
2.1. The Hausdorff extension theorem
The following celebrated theorem was first proven by Hausdorff [14] (cf. [15], [6], [2], and [28]). The latter part on complete metrics was added by Bacon [1].
Theorem 2.1.
Let $X$ be a metrizable space, and let $A$ be a closed subset of $X$. Then for every $d\in\mathrm{Met}(A)$ there exists $D\in\mathrm{Met}(X)$ with $D|_{A^{2}}=d$. Moreover, if $X$ is completely metrizable, and if $d\in\mathrm{Met}(A)$ is a complete metric on $A$, then we can choose $D\in\mathrm{Met}(X)$ as a complete metric on $X$.
2.2. The Michael continuous selection theorem
Let $V$ be a Banach space. We denote by $\operatorname{\mathcal{CC}}(V)$ the set of all non-empty closed convex subsets of $V$. For a topological space $X$ we say that a map $\phi\colon X\to\operatorname{\mathcal{CC}}(V)$ is lower semi-continuous if for every open subset $O$ of $V$ the set $\{\,x\in X\mid\phi(x)\cap O\neq\emptyset\,\}$ is open in $X$.
The following theorem is known as one of the Michael continuous selection theorems proven in [23].
Theorem 2.2.
Let $X$ be a paracompact space, and $A$ a closed subsets of $X$, $V$ a Banach space, and $\phi\colon X\to\operatorname{\mathcal{CC}}(V)$ be a lower semi-continuous map. If a continuous map $f\colon A\to B$ satisfies $f(x)\in\phi(x)$ for all $x\in A$, then there exists a continuous map $F\colon X\to V$ with $F|_{A}=f$ such that for every $x\in X$ we have $F(x)\in\phi(x)$.
By using linear structure, we have the following:
Proposition 2.3.
Let $V$ be a Banach space, and $x,y\in V$. Then for every $r\in(0,\infty)$ we have $\mathcal{H}(B(x,r),B(y,r))=\|x-y\|_{V}$, where $\mathcal{H}$ is the Hausdorff distance induced from the norm $\|\cdot\|_{V}$ of $V$.
Corollary 2.4.
Let $X$ be a topological space, $V$ be a Banach space, and let $H\colon X\to V$ be a continuous map and $r\in(0,\infty)$. Then a map $\phi\colon X\to\operatorname{\mathcal{CC}}(V)$ defined by $\phi(x)=B(H(x),r)$ is lower semi-continuous.
Proof.
For every open subset $O$ of $V$, and for every point $a\in X$ with $\phi(a)\cap O\neq\emptyset$, choose $u\in\phi(a)\cap O$ and $l\in(0,\infty)$ with $U(u,l)\subset O$. By Proposition 2.3, we can take $\delta\in(0,\infty)$ such that for every $x\in U(a,\delta)$ we have $\mathcal{H}(\phi(x),\phi(a))=\|H(x)-H(a)\|_{V}<l$. Then we have $\phi(x)\cap U(u,l)\neq\emptyset$, and hence $\phi(x)\cap O\neq\emptyset$. Therefore the set $\{\,x\in X\mid\phi(x)\cap O\neq\emptyset\,\}$ is open in $X$. ∎
The following theorem is known as the Stone theorem proven in [27].
Theorem 2.5.
All metrizable spaces are paracompact.
By this theorem, we can apply Theorem 2.2 to all metrizable space.
2.3. The Kuratowski embedding theorem
For a metric space $(X,d)$, we denote by $C_{b}(X)$ the Banach space of all bounded continuous functions on $X$ equipped with the supremum norm. For $x\in X$, we denote by $d_{x}$ the function from $X$ to $\mathbb{R}$ defined by $d_{x}(p)=d(x,p)$. The following theorem, which states that every metric space is isometrically embeddable into Banach spaces, is known as the Kuratowski embedding theorem. The proof can be seen in, for example, [11, Theorem 4.3.14].
Theorem 2.6.
Let $(X,d)$ be a metric space. Take $o\in X$. Then the map $K:X\to C_{b}(X)$ defined by $K(x)=d_{x}-d_{o}$ is an isometric embedding. Moreover, if $(X,d)$ is bounded, the map $L\colon X\to C_{b}(X)$ defined by $L(x)=d_{x}$ is an isometric embedding.
2.4. Baire spaces
A topological space $X$ is said to be Baire if the intersections of countable dense open subsets of $X$ are dense in $X$.
The following is known as the Baire category theorem.
Theorem 2.7.
Every completely metrizable space is a Baire space.
Since $G_{\delta}$ subset of completely metrizable space is completely metrizable (see, e.g. [31, Theorem 24.12]), we obtain the following:
Lemma 2.8.
Every $G_{\delta}$ subset of a completely metrizable space is a Baire space.
3. An interpolation Theorem of metrics
In this section, we prove Theorem 1.1.
3.1. Amalgamation lemmas
Let $X$ and $Y$ be sets, and let $\tau\colon X\to Y$ be a bijection. For a metric $d$ on $Y$, we denote by $\tau^{*}d$ the metric on $X$ defined by $(\tau^{*}d)(x,y)=d(\tau(x),\tau(y))$. In this case, the map $\tau$ is an isometry between $(X,\tau^{*}d)$ and $(Y,d)$.
The following proposition can be considered as a specific case of the realization of the Gromov–Hausdorff distance of two metric spaces.
Proposition 3.1.
Let $X$ be a metrizable space. For $r\in(0,\infty)$, let $d,e\in\mathrm{Met}(X)$ with $\mathcal{D}_{X}(d,e)\leq r$. Put $X_{0}=X$, and let $X_{1}$ be a set with $\operatorname{card}(X_{1})=\operatorname{card}(X_{0})$ and $X_{0}\cap X_{1}=\emptyset$. Take a bijection $\tau:X_{0}\to X_{1}$. Then there exists a metric $h\in\mathrm{Met}(X_{0}\sqcup X_{1})$ such that
  1. (1)
    $h|_{(X_{0})^{2}}=d$;
  2. (2)
    $h|_{(X_{1})^{2}}=(\tau^{-1})^{*}e$;
  3. (3)
    for every $x\in X_{0}$ we have $h(x,\tau(x))=r/2$.
Proof.
We define a symmetric function $h:(X_{0}\sqcup X_{1})^{2}\to[0,\infty)$ by
$h(x,y)=\begin{cases}d(x,y)&\text{if $x,y\in X_{0}$;}\\ e(x,y)&\text{if $x,y\in X_{1}$;}\\ \inf_{a\in X_{0}}(d(x,a)+r/2+e(\tau(a),y))&\text{if $(x,y)\in X_{0}\times X_{1}$. }\end{cases}$
Then $h$ is a metric as desired. For more details, we refer the readers to [8, Theorem 7.3.25]. ∎
The following proposition is known as an amalgamation of metrics.
Proposition 3.2.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces, and let $Z=X\cap Y$. If $Z\neq\emptyset$ and $d_{X}|_{Z^{2}}=d_{Y}|_{Z^{2}}$, then there exists a metric $h$ on $X\cup Y$ such that
  1. (1)
    $h|_{X^{2}}=d_{X}$;
  2. (2)
    $h|_{Y^{2}}=d_{Y}$.
Proof.
We define a symmetric function $h:(X\cup Y)^{2}\to[0,\infty)$ by
$h(x,y)=\begin{cases}d_{X}(x,y)&\text{if $x,y\in X$;}\\ d_{Y}(x,y)&\text{if $x,y\in Y$;}\\ \inf_{z\in Z}(d_{X}(x,z)+d_{Y}(z,y))&\text{if $(x,y)\in X\times Y$. }\end{cases}$
Since $d_{X}|_{Z^{2}}=d_{Y}|_{Z^{2}}$, the function $h$ is well-defined, and it becomes a desired metric. The detailed proof can be seen in, for example, [7] (cf. [30] and [10]). ∎
For a mutually disjoint family $\{T_{i}\}_{i\in I}$ of topological spaces, we consider that the space $\coprod_{i\in I}T_{i}$ is always equipped with the direct sum topology.
Proposition 3.3.
Let $X$ be a metrizable space, and $\{A_{i}\}_{i\in I}$ be a discrete family of closed subsets of $X$. Take $d\in\mathrm{Met}(X)$, and let $\{e_{i}\}_{i\in I}$ be a family of metrics with $e_{i}\in\mathrm{Met}(A_{i})$. Put $\eta=\sup_{i\in I}\mathcal{D}_{A_{i}}(e_{A_{i}},d|_{(A_{i})^{2}})$ and assume that $\eta<\infty$. Let $\{B_{i}\}_{i\in I}$ be a mutually disjoint family of sets such that for all $i\in I$ we have $X\cap B_{i}=\emptyset$. Take $\tau\colon\coprod_{i\in I}A_{i}\to\coprod_{i\in I}B_{i}$ be a bijection such that for each $i\in I$ the map $\tau_{i}=\tau|_{A_{i}}$ is a bijection between $A_{i}$ and $B_{i}$. Then there exists a metric $h$ on $X\sqcup\coprod_{i\in I}B_{i}$ such that
  1. (1)
    for every $i\in I$ we have $h|_{(B_{i})^{2}}=(\tau_{i}^{-1})^{*}e_{i}$;
  2. (2)
    $h|_{X^{2}}=d$;
  3. (3)
    for every $x\in\coprod_{i\in I}A_{i}$ we have $h(x,\tau(x))=\eta/2$.
Proof.
for every $i\in I$, Proposition 3.1 enables us to take a metric $l_{i}\in\mathrm{Met}(A_{i}\sqcup B_{i})$ such that
  1. (a1)
    $l_{i}|_{A_{i}^{2}}=d|_{A_{i}^{2}}$;
  2. (a2)
    $l_{i}|_{B_{i}^{2}}=(\tau_{i}^{-1})^{*}e_{i}$;
  3. (a3)
    for all $x\in A_{i}$ we have $l_{i}(x,\tau(x))=\eta/2$.
From now on, we construct a metric mentioned in the proposition by transfinite recursion. We may assume that the index set $I$ is a cardinal $\kappa$. For each $\gamma<\theta$, we put $Z_{\gamma}=X\sqcup\coprod_{\alpha<\gamma}B_{\alpha}$.
Fix $\theta\leq\kappa$, and assume that we already constructed a metric $\{h_{\alpha}\}_{\alpha<\theta}$ such that
  1. (b1)
    If $\alpha<\beta<\theta$, then we have $h_{\beta}|Z_{\alpha}=h_{\alpha}$.
  2. (b2)
    If $\alpha<\beta<\theta$, then we have $h_{\beta}|_{(A_{\alpha}\cup B_{\alpha})^{2}}=l_{\alpha}$.
Under this assumptions, we shall make $h_{\theta}$. We divide the construction into two cases.
If $\theta=\gamma+1$ for some $\gamma$, then we apply Proposition 3.2 to $h_{\gamma}$ and $l_{\gamma}$, and we then obtain $h_{\theta}$ satisfying (b1) and (b2) for $\theta+1$.
If $\theta$ is a limit cardinal, then we define a metric $h_{\theta}$ on $Z_{\theta}$ by $h_{\theta}(x,y)=h_{\alpha}(x,y)$, where $\alpha<\theta$ satisfies that $x,y\in Z_{\alpha}$. The existence of $\alpha$ is guaranteed by $Z_{\theta}=\bigcup_{\alpha<\theta}Z_{\alpha}$. By (b1), the metric $h_{\theta}$ is well-defined. Similarly, we see that $h_{\theta}$ satisfies (b1) and (b2) for $\theta+1$.
Put $h=h_{\kappa}$. From the assumptions (b1) and (b2) for $\kappa$, and from $Z_{\kappa}=X\sqcup\coprod_{i\in I}B_{i}$, it follows that $h$ is as required. ∎
3.2. Proof of Theorem 1.1
To show Theorem 1.1, we first prove the case where the index set is a singleton.
Theorem 3.4.
Let $X$ be a metrizable space, and let $A$ be a closed subset of $X$. Then for every $d\in\mathrm{Met}(X)$, and for every $e\in\mathrm{Met}(A)$, there exists a metric $m\in\mathrm{Met}(X)$ such that:
  1. (1)
    $m|_{A^{2}}=e$;
  2. (2)
    $\mathcal{D}_{X}(m,d)=\mathcal{D}_{A}(e,d|_{A^{2}})$.
Moreover, if $X$ is completely metrizable, and if $e\in\mathrm{Met}(A)$ is a complete metric, then we can choose $m\in\mathrm{Met}(X)$ as a complete metric.
Proof.
Put $\eta=\mathcal{D}_{A}(e,d|_{A^{2}})$. If $\eta=\infty$, then the theorem follows from the Hausdorff extension theorem 2.1.
We may assume $\eta<\infty$. Let $B$ be a set such that $X\cap B=\emptyset$ and $\operatorname{card}(B)=\operatorname{card}(A)$, and let $\tau\colon A\to B$ be a bijection. Put $Z=X\sqcup B$. By Proposition 3.3, we find a metric $h$ on $Z$ such that
  1. (H1)
    we have $h|_{B^{2}}=(\tau^{-1})^{*}e$;
  2. (H2)
    $h|_{X^{2}}=d$;
  3. (H3)
    for every $x\in A$ we have $h(x,\tau(x))=\eta/2$.
We can take an isometric embedding $H\colon Z\to Y$ from $(Z,h)$ into a Banach space $(Y,\|\cdot\|_{Y})$ (see e.g., the Kuratowski embedding theorem 2.6). Define a map $\phi\colon Z\to\operatorname{\mathcal{CC}}(Y)$ by $\phi(x)=B(H(x),\eta/2)$. By Corollary 2.4, the map $\phi$ is lower semi-continuous. We define a map $f\colon A\to Y$ by $f(x)=H(\tau(x))$. Then $f$ is continuous. By the property (H3) of $h$, for every $x\in A$ we have $f(x)\in\phi(x)$. Due to the Stone theorem 2.5, the space $X$ is paracompact. Thus we can apply the Michael selection theorem 2.2 to the map $f$, and hence we obtain a continuous map $F\colon X\to Y$ such that $F|_{A}=f$ and for every $x\in A$ we have $F(x)\in\phi(x)$. Note that $F(x)\in\phi(x)$ means that
$\|F(x)-H(x)\|_{Y}\leq\eta/2.$
Define a map $l\colon X\times X\to[0,\infty)$ by $l(x,y)=\min\{e(x,y),\eta/2\}$. Note that $l\in\mathrm{Met}(X)$. We consider that the product Banach space $Y\times C_{b}(X)$ is equipped with the max norm defined by $\|(x,y)\|=\max\{\|x\|_{Y},\|y\|_{C_{b}(X)}\}$.
Define a map $E\colon X\to Y\times C_{b}(X)$ by
$E(x)=(F(x),l_{x}),$
where $l_{x}$ is a bounded function on $X$ defined by $l_{x}(p)=l(x,p)$. By the Kuratowski embedding theorem 2.6, the map $L\colon X\to C_{b}(X)$ defined by $L(x)=l_{x}$ is an isometric embedding. Therefore $E$ is a topological embedding. We also define a map $K\colon X\to Y\times C_{b}(X)$ by
$K(x)=(H(x),0).$
Then, by the definition of the norm of $Y\times C_{b}(X)$, the map $K$ from $(X,d)$ to $(Y\times C_{b}(X),\|\cdot\|)$ is an isometry. Since for every $x\in X$ we have $\|F(x)-H(x)\|_{Y}\leq\eta/2$ and $\|l_{x}\|_{C_{b}(X)}\leq\eta/2$, we obtain
(3.1) $\|E(x)-K(x)\|=\max\{\|F(x)-H(x)\|_{Y},\|l_{x}\|_{C_{b}(X)}\}\leq\eta/2.$
Define a function $m:X^{2}\to[0,\infty)$ by $m(x,y)=\|E(x)-E(y)\|$, then $m$ is a metric on $X$. Since $E$ is a topological embedding, we see that $m\in\mathrm{Met}(X)$. For every pair $x,y\in A$, we have $\|F(x)-F(y)\|_{Y}=e(x,y)$ and
$\|l_{x}-l_{y}\|_{C_{b}(X)}=l(x,y)\leq e(x,y);$
thus we obtain
$\|E(x)-E(y)\|=\max\{\|F(x)-F(y)\|_{Y},\|l_{x}-l_{y}\|_{C_{b}(X)}\}=e(x,y),$
and hence $m|_{A^{2}}=e$. Moreover, we have $\eta\leq\mathcal{D}_{X}(m,d)$. We also obtain the opposite inequality $\mathcal{D}_{X}(m,d)\leq\eta$; indeed, the inequality (3.1) shows that for every pair $x,y\in X$, we have
[…]
$\leq\|E(y)-K(y)\|+\|E(x)-K(x)\|\leq\eta/2+\eta/2=\eta.$
Therefore we conclude that $\mathcal{D}_{X}(m,d)=\eta$. This completes the proof of the former part of Theorem 1.1.
By the latter part of the Hausdorff theorem 2.1, we can choose $l$ as a complete metric. Then $m$ become a complete metric. This shows the latter part of Theorem 1.1. ∎
Proof of Theorem 1.1.
Let $X$ be a metrizable space, and let $\{A_{i}\}_{i\in I}$ be a discrete family of closed subsets of $X$. Put $A=\bigcup_{i\in I}A_{i}$. Since $\{A_{i}\}_{i\in I}$ is a discrete family, the union set $A=\bigcup_{i\in I}A_{i}$ is also closed in $X$. We use the same notation of Proposition 3.3. We define $w=\tau^{*}h\in\mathrm{Met}(A)$. note that $w|_{A_{i}^{2}}=e_{i}$ for every $i\in I$. Let us check that $\mathcal{D}_{A}(d|_{A^{2}},w)=\eta$. In fact, for every pair $x,y\in A$, we have
$|d(x,y)-w(x,y)|$ $\leq|h(x,y)-h(x,\tau(y))|+|h(x,\tau(y))-h(\tau(x),\tau(y))|$
$\leq h(y,\tau(y))+h(x,\tau(x))\leq\eta/2+\eta/2=\eta.$
Hence $\mathcal{D}_{A}(d|_{A^{2}},w)=\eta$. Applying Theorem 3.4 to $d$ and $w$, we obtain $m\in\mathrm{Met}(X)$ such that $\mathcal{D}_{X}(d,m)=\eta$ and $m|_{A^{2}}=w$. This finishes the proof of Theorem 1.1. ∎
4. Transmissible properties
In this section we discuss transmissible properties, and prove Theorem 1.2. We also show that various properties in metric geometry are transmissible properties.
4.1. Proof of Theorem 1.2
By the condition (TP2) in Definition 1.1, we obtain the following:
Lemma 4.1.
Let $\mathfrak{G}$ be a transmissible parameter. If a metric space $(X,d)$ satisfies the $\mathfrak{G}$-transmissible property, then so does every metric subspace of $(X,d)$.
By the virtue of Lemma 4.1, we use the word “transmissible”.
Corollary 4.2.
Let $\mathfrak{G}$ be a transmissible parameter, and let $(X,d)$ be a metric space. If there exists a metric subspace of $(X,d)$ satisfies the anti-$\mathfrak{G}$-transmissible property, then so does $(X,d)$.
Let $X$ be a metrizable space, and let $\mathfrak{G}=(Q,P,F,G,Z,\phi)$ be a transmissible parameter. For $q\in Q$, for $a\in\operatorname{Seq}(G(q),X)$ and for $z\in Z$, we denote by $S(X,\mathfrak{G},q,a,z)$ the set of all $d\in\mathrm{Met}(X)$ such that $\phi^{q,X}(a,z,d)\in X\setminus F(q)$. We also denote by $S(X,\mathfrak{G})$ the set of all $d\in\mathrm{Met}(X)$ such that $(X,d)$ satisfies the anti-$\mathfrak{G}$-transmissible property.
Proposition 4.3.
Let $X$ be a metrizable space, $\mathfrak{G}=(Q,P,F,G,Z,\phi)$ a transmissible parameter. Then for every $q\in Q$, for every $a\in\operatorname{Seq}(G(q),X)$, and for every $z\in Z$, the set $S(X,\mathfrak{G},q,a,z)$ is open in $\mathrm{Met}(X)$.
Proof.
Fix $q\in Q$, $a\in\operatorname{Seq}(G(q),X)$ and $z\in Z$. Since the map $\phi^{q,X}(a,z)\colon\mathrm{Met}(X)\to P$ is continuous and since $X\setminus F(q)$ is open in $P$, the set $S(X,\mathfrak{G},q,a,z)$ is open in $\mathrm{Met}(X)$. ∎
Corollary 4.4.
Let $X$ be a metrizable space, $\mathfrak{G}=(Q,P,F,G,Z,\phi)$ a transmissible parameter. Then the set $S(X,\mathfrak{G})$ is $G_{\delta}$ in $\mathrm{Met}(X)$. Moreover, if the set $Q$ is finite, then $S(X,\mathfrak{G})$ is open in $\mathrm{Met}(X)$.
Proof.
By the definitions of $S(X,\mathfrak{G})$ and $S(X,\mathfrak{G},q,a,z)$, we have
$S(X,\mathfrak{G})=\bigcap_{q\in Q}\bigcup_{a\in\operatorname{Seq}(G(q),X)}\bigcup_{z\in Z}S(X,\mathfrak{G},q,a,z).$
This equality and Proposition 4.3 prove the lemma. ∎
We say that a topological space is an $(\omega_{0}+1)$-space if it is homeomorphic to the one-point compactification of the countable discrete topological space. Namely, it is homeomorphic to the ordinal $\omega_{0}+1$.
Lemma 4.5.
Let $\mathfrak{G}$ be a singular transmissible parameter. Then there exists an $(\omega_{0}+1)$-metric space with arbitrary small diameter satisfying the anti-$\mathfrak{G}$-transmissible property.
Proof.
Let $\mathfrak{G}=(Q,P,F,G,Z,\phi)$. Fix $\epsilon\in(0,\infty)$. By the singularity of $\mathfrak{G}$, there exists a sequence $\{(R_{i},d_{i})\}_{i\in\mathbb{N}}$ of finite metric spaces such that for each $i\in\mathbb{N}$ there exist $q_{i}\in Q$ and $z_{i}\in Z$ satisfying
  1. (R1)
    $\delta_{d_{i}}(R_{i})\leq\epsilon\cdot 2^{-i}$;
  2. (R2)
    $\operatorname{card}(R_{i})\in G(q_{i})$;
  3. (R3)
    $\phi^{q_{i},R_{i}}(R_{i},z_{i},d_{i})\in X\setminus F(q_{i})$.
Put
$L=\{\infty\}\sqcup\coprod_{i\in\mathbb{N}}R_{i},$
and define a metric $d_{L}$ on $L$ by
[…]
On account of (R1), $d_{L}$ is actually a metric (compare with [18, Definition 3.3]). We also observe that the space $(L,d_{L})$ is an $(\omega_{0}+1)$-metric space with $\delta_{d_{L}}(L)\leq\epsilon$. By the properties (R2) and (R3) of $\{(R_{i},d_{i})\}_{i\in\mathbb{N}}$, the metric space $(L,d_{L})$ satisfies the anti-$\mathfrak{G}$-transmissible property. ∎
Remark 4.1.
It is also true that a transmissible parameter $\mathfrak{G}$ is singular if and only if there exists an $(\omega_{0}+1)$-metric space with arbitrary small diameter satisfying the anti-$\mathfrak{G}$-transmissible property.
Let $\mathfrak{G}$ be a transmissible parameter. For a non-discrete metrizable space $X$, and for an $(\omega_{0}+1)$-subspace $R$ of $X$, we denote by $T(X,R,\mathfrak{G})$ the set of all $d\in\mathrm{Met}(X)$ for which $(R,d|_{R^{2}})$ satisfies the anti-$\mathfrak{G}$-transmissible property.
As a consequence of Theorem 3.4, we obtain the following:
Proposition 4.6.
Let $\mathfrak{G}=(Q,P,F,G,Z,\phi)$ be a singular transmissible parameter. Then for every non-discrete metrizable space $X$, and for every $(\omega_{0}+1)$-subspace $R$ of $X$, the set $T(X,R,\mathfrak{G})$ is dense in $\mathrm{Met}(X)$.
Proof.
Fix $d\in\mathrm{Met}(X)$ and $\epsilon\in(0,\infty)$. From the singularity of $\mathfrak{G}$, by Lemma 4.5, it follows that there exists an $(\omega_{0}+1)$-metric space $(L,e)$ satisfying the anti-$\mathfrak{G}$-transmissible property and $\delta_{e}(L)<\epsilon/2$. Since $R$ is an $(\omega_{0}+1)$-space, there exists an $(\omega_{0}+1)$ subspace $S$ of $R$ with $\delta_{d}(S)<\epsilon/2$. Let $\tau\colon S\to L$ be a homeomorphism. By the definitions of $S$ and $e$, we have $\mathcal{D}_{S}(d|_{S^{2}},\tau^{*}e)<\epsilon$. Theorem 3.4 guarantees the existence of a metric $m\in\mathrm{Met}(X)$ such that $m|_{S^{2}}=\tau^{*}e$ and $\mathcal{D}_{X}(m,d)<\epsilon$. Due to Corollary 4.2, the metric space $(R,m|_{R^{2}})$ satisfies the anti-$\mathfrak{G}$-transmissible property. Since $d$ and $\epsilon$ are arbitrary, we conclude that $T(X,R,\mathfrak{G})$ is dense in $\mathrm{Met}(X)$. ∎
Proof of Theorem 1.2.
Let $X$ be a non-discrete metrizable space, and let $\mathfrak{G}$ be a singular transmissible parameter. Since $X$ is non-discrete, there exists an $(\omega_{0}+1)$-subspace $R$ of $X$. By the definitions, we have
$T(X,R,\mathfrak{G})\subset S(X,\mathfrak{G}).$
By Proposition 4.6 and Corollary 4.4, the set $S(X,\mathfrak{G})$ is dense $G_{\delta}$ in $\mathrm{Met}(X)$. This finishes the proof. ∎
For a complete metrizable space $X$, we denote by $\operatorname{Comp}(X)$ the set of all complete metrics in $\mathrm{Met}(X)$. From the latter part of Theorem 3.4, we deduce the following:
Theorem 4.7.
Let $\mathfrak{G}$ be a singular transmissible parameter. For every non-discrete completely metrizable space $X$, the set of all $d\in\operatorname{Comp}(X)$ for which $(X,d)$ satisfies the anti-$\mathfrak{G}$-transmissible property is dense $G_{\delta}$ in $\operatorname{Comp}(X)$.
4.2. The doubling property and the uniform disconnectedness
For a metic space $(X,d)$ and for a subset $A$ of $X$, we set
$\alpha_{d}(A)=\inf\{\,d(x,y)\mid\text{$x,y\in A$ and $x\neq y$}\,\}.$
A metric space $(X,d)$ is said to be doubling if there exist $C\in(0,\infty)$ and $\lambda\in(0,\infty)$ such that for every finite subset $A$ of $X$ we have
$\operatorname{card}(A)\leq C\left(\frac{\delta_{d}(A)}{\alpha_{d}(A)}\right)^{\lambda}.$
Note that $(X,d)$ is doubling if and only if $(X,d)$ has finite Assouad dimension (see e.g., [16, Section 10]).
By the definitions of the topology of $\mathrm{Met}(X)$, $\alpha_{d}$ and $\delta_{d}$, we obtain:
Lemma 4.8.
Let $X$ be a metrizable space. Fix a finite subset $A$ of $X$. Then maps $\alpha_{*,A},\delta_{*,A}\colon\mathrm{Met}(X)\to\mathbb{R}$ defined by $\alpha_{*,A}(d)=\alpha_{d}(A)$ and $\delta_{*,A}(d)=\delta_{d}(A)$ is continuous.
Proposition 4.9.
The doubling property on metric spaces is a transmissible property with a singular transmissible parameter.
Proof.
Define a map $F_{D}\colon(\mathbb{Q}_{>0})^{2}\to\mathcal{F}((\mathbb{R}_{>0})^{2})$ by
$F_{D}((q_{1},q_{2}))=\{\,(x,y)\in(\mathbb{R}_{>0})^{2}\mid x\leq q_{1}\cdot y^{q_{2}}\,\},$
and define a constant map $G_{D}\colon(\mathbb{Q}_{>0})^{2}\to\mathcal{P}(\mathbb{N})^{*}$ by $G_{D}(q)=[2,\infty)$. Put $Z_{D}=\{1\}$. For each metrizable space $X$, and for each $q\in(\mathbb{Q}_{>0})^{2}$, define a map $\phi_{D}^{q,X}\colon\operatorname{Seq}(G(q),X)\times Z_{D}\times\mathrm{Met}(X)\to(\mathbb{R}_{>0})^{2}$ by
$\phi_{D}^{q,X}(\{a_{i}\}_{i=1}^{N},1,d)=\left(N,\frac{\delta_{d}(\{\,a_{i}\mid i\in\{1,\dots,N\}\,\})}{\alpha_{d}(\{\,a_{i}\mid i\in\{1,\dots,N\}\,\})}\right).$
Let $\mathfrak{DB}=((\mathbb{Q}_{>0})^{2},(\mathbb{R}_{>0})^{2},F_{D},G_{D},\{1\},\phi_{D})$. Then $\mathfrak{DB}$ satisfies the condition (TP2) in Definition 1.1. By the Lemma 4.8, we see that $\mathfrak{DB}$ satisfies the condition (TP1). Hence $\mathfrak{DB}$ is a transmissible parameter. The $\mathfrak{DB}$-transmissible property is equivalent to the doubling property. We next prove that $\mathfrak{DB}$ is singular. For $q=(q_{1},q_{2})\in(\mathbb{Q}_{>0})^{2}$ and for $\epsilon\in(0,\infty)$, we denote by $(R_{q},d_{q})$ a finite metric space with $\operatorname{card}(R_{q})>q_{1}+1$ and $d_{q}(x,y)=\epsilon$ whenever $x\neq y$. Then $\delta_{d_{q}}(R_{q})=\epsilon$, and
$\phi^{q,R_{q}}(R_{q},1,d_{q})=(\operatorname{card}(R_{q}),1)\not\in F_{D}(q).$
This implies the proposition. ∎
Remark 4.2.
Let $\mathbb{Z}$ be the set of all integers with discrete topology, and let $h\in\mathrm{Met}(\mathbb{Z})$ be the relative metric on $\mathbb{Z}$ induced from the Euclidean metric on $\mathbb{R}$. Then $h$ has a neighborhood $U$ in $\mathrm{Met}(X)$ such that for every $d\in U$ the space $(\mathbb{Z},d)$ is doubling.
A metric space $(X,d)$ is said to be uniformly disconnected if there exists $\delta\in(0,1)$ such that if a finite sequence $\{z_{i}\}_{i=1}^{N}$ in $X$ satisfies $\max_{i}d(z_{i},z_{i+1})<\delta d(z_{1},z_{N})$, then we have $N=1$. Namely, the inequality $N>1$ implies $\max_{i}d(z_{i},z_{i+1})\geq\delta d(z_{1},z_{N})$. Note that a metric space is uniformly disconnected if and only if it is bi-Lipschitz to an ultrametric space (see e.g., [22, Lemma 5.1.10]).
By the definition of the topology of $\mathrm{Met}(X)$, we obtain:
Lemma 4.10.
Let $X$ be a metrizable space. Fix two points $a$, $b$ in $X$. Then a map $f\colon\mathrm{Met}(X)\to\mathbb{R}$ defined by $f(d)=d(a,b)$ is continuous.
Proposition 4.11.
The uniform disconnectedness on metic spaces is a transmissible property with a singular parameter.
Proof.
Define a map $F_{UD}\colon\mathbb{Q}\cap(0,1)\to\mathcal{F}((\mathbb{R}_{\geq 0})^{2})$ by
$F_{UD}(q)=\{\,(x,y)\in(\mathbb{R}_{\geq 0})^{2}\mid x\geq qy\,\},$
and define a constant map $G_{UD}\colon\mathbb{Q}\cap(0,1)\to\mathcal{P}^{*}(\mathbb{N})$ by $G_{UD}(q)=[2,\infty)$. Put $Z_{UD}=\{1\}$. For each metrizable space $X$, and for each $q\in\mathbb{Q}\cap(0,1)$, define a map $\phi_{UD}^{q,X}\colon\operatorname{Seq}(G_{UD}(q),X)\times Z_{UD}\times\mathrm{Met}(X)\to(\mathbb{R}_{\geq 0})^{2}$ by
$\phi_{UD}^{q,X}(\{a_{i}\}_{i=1}^{N},1,d)=\left(\max_{1\leq i\leq N-1}d(a_{i},a_{i+1}),d(a_{1},a_{N})\right).$
Let $\mathfrak{UD}=(\mathbb{Q}\cap(0,1),(\mathbb{R}_{>0})^{2},F_{UD},G_{UD},\{1\},\phi_{UD})$. Then $\mathfrak{UD}$ satisfies the conditions (TP2) in Definition 1.1. By Lemma 4.10, we see that $\mathfrak{UD}$ satisfies the condition (TP1). Hence $\mathfrak{UD}$ is a transmissible parameter, and the $\mathfrak{UD}$-transmissible property is equivalent to the uniform disconnectedness. We next prove that $\mathfrak{UD}$ is singular. For every $q\in\mathbb{Q}\cap(0,1)$, take $n\in\mathbb{N}$ with $1/n<q$. Put
$R_{q}=\{\,\epsilon\cdot i/n\mid i\in\mathbb{Z}\cap[0,n]\,\},$
and let $d_{q}$ be the relative metric on $R_{q}$ induced from the Euclidean metric. Then $\delta_{d_{q}}(R_{q})=\epsilon$, and
$\phi_{UD}^{q,R_{q}}(\{a_{i}\}_{i=1}^{N},1,d_{q})=(\epsilon/n,\epsilon)\not\in U(q).$
This leads to the proposition. ∎
Remark 4.3.
Let $C$ be a countable discrete space, and let $h\in\mathrm{Met}(C)$ be a metric such that $h(x,y)=1$ whenever $x\neq y$. Then $h$ has a neighborhood $U$ in $\mathrm{Met}(C)$ such that for every $d\in U$ the space $(C,d)$ is uniformly disconnected.
From Propositions 4.9 and 4.11 and from Theorem 1.2, we deduce the next observations on dense $G_{\delta}$ subsets of moduli spaces of metrics.
Theorem 4.12.
For every non-discrete metrizable space $X$, the following two sets are dense $G_{\delta}$ in $\mathrm{Met}(X)$.
  1. (1)
    the set of all metrics $d\in\mathrm{Met}(X)$ for which $(X,d)$ is not doubling;
  2. (2)
    the set of all metrics $d\in\mathrm{Met}(X)$ for which $(X,d)$ is not uniformly disconnected.
4.3. Rich pseudo-cones
Let $(X,d)$ be a metric space. Let $\{A_{i}\}_{i\in\mathbb{N}}$ be a sequence of subsets of $X$, and let $\{u_{i}\}_{i\in\mathbb{N}}$ be a sequence in $(0,\infty)$. We say that a metric space $(P,d_{P})$ is a pseudo-cone of $X$ approximated by $(\{A_{i}\}_{i\in\mathbb{N}},\{u_{i}\}_{i\in\mathbb{N}})$ if
$\lim_{i\to\infty}\mathcal{GH}((A_{i},u_{i}\cdot d|_{A_{i}^{2}}),(P,d_{P}))=0$
(see [19]), where $\mathcal{GH}$ is the Gromov–Hausdorff distance (see [8]). For a metric space $(X,d)$, we denote by $\operatorname{\mathscr{P}}(X,d)$ the class of all pseudo-cones of $(X,d)$. Let $\mathscr{F}$ be the class of all finite metric spaces whose distances are rational numbers. We denote by $\mathscr{G}$ the quotient class of $\mathscr{F}$ divided by the isometric equivalence. Note that $\mathscr{G}$ is countable.
We say that a metric space $(X,d)$ has rich pseudo-cones if $\mathscr{F}$ is contained in $\operatorname{\mathscr{P}}(X,d)$.
Proposition 4.13.
The rich pseudo-cones property on metric spaces is an anti-transmissible property with a singular transmissible parameter.
Proof.
Let $\{(L_{n},d_{n})\}_{n\in\mathbb{N}}$ be a complete representation system of $\mathscr{G}$. Let $L_{n}=\{f_{n,l}\}_{l=1}^{\operatorname{card}(F_{i})}$. Define a function $F_{R}\colon\mathbb{N}^{2}\to\mathcal{F}(\mathbb{R})$ by
$F_{R}(n,m)=\{\,y\in\mathbb{R}\mid y\geq 2^{-m}\,\},$
and define a map $G_{R}\colon\mathbb{N}^{2}\to\mathcal{P}^{*}(\mathbb{N})$ by $G_{R}(n,m)=\{\operatorname{card}(L_{n})\}$.
For each $k=(n,m)\in\mathbb{N}^{2}$, for each metrizable space $X$, for each $\{a_{i}\}_{i=1}^{M}\in\operatorname{Seq}(G_{R}(k),X)$, and for all $i,j\in\{1,\dots,M\}$ we define a function $r_{i,j}^{k}(\{a_{i}\}_{i=1}^{M})\colon(0,\infty)\times\mathrm{Met}(X)\to\mathbb{R}$ by
$r_{i,j}^{k}(\{a_{i}\}_{i=1}^{M})(z,d)=|z^{-1}d(a_{i},a_{j})-d_{n}(f_{n,i},f_{n,j})|$
if $i,j\in\{1,\dots,M\}$; otherwise, we define $r_{i,j}^{k}(\{a_{i}\}_{i=1}^{M})(z,d)=0$. By Lemma 4.10, the map $r_{i,j}^{k}(\{a_{i}\}_{i=1}^{M})$ is continuous.
Define a map $\phi_{R}^{k,X}\colon\operatorname{Seq}(G(k),X)\times(0,\infty)\times\mathrm{Met}(X)\to\mathbb{R}$ by
$\phi_{R}^{k,X}(\{a_{i}\}_{i=1}^{M},z,d)=\max_{i,j\in\{1,\dots,M\}}r_{i,j}^{k}(\{a_{i}\}_{i=1}^{M})(z,d).$
Let $\mathfrak{R}=(\mathbb{N}^{2},\mathbb{R},F_{R},G_{R},(0,\infty),\phi_{r})$. Then $\mathfrak{R}$ satisfies the conditions (TP1) and (TP2) in Definition 1.1, and hence it is a transmissible parameter.
For a metric space $(X,d)$, the anti-$\mathfrak{R}$-transmissible property means that for every $n\in\mathbb{N}$, and for every $m\in\mathbb{N}$, there exist a finite subspace $A=\{a_{i}\}_{i=1}^{\operatorname{card}(L_{n})}$ of $X$ and a positive number $z\in(0,\infty)$ such that for all $i,j\in\{1,\dots,\operatorname{card}(L_{n})\}$ we have
$|z^{-1}d(a_{i},a_{j})-d_{n}(f_{n,i},f_{n,j})|<2^{-m};$
in particular, $\mathcal{GH}((A,z^{-1}d|_{A^{2}}),(L_{n},d_{n}))<2^{-(m+1)}$. Thus $\mathscr{F}$ is contained in $\operatorname{\mathscr{P}}(X,d)$. This implies that $(X,d)$ has rich pseudo-cones. We next prove the opposite. If $(X,d)$ has rich pseudo-cones, then for every $(W,d_{W})\in\mathscr{F}$, and for every $\epsilon\in(0,\infty)$, there exist a positive number $z\in(0,\infty)$ and a subset $A$ of $X$ with $\operatorname{card}(A)=\operatorname{card}(W)$ such that $\mathcal{GH}((A,z^{-1}d|_{A^{2}}),(W,d_{W}))<\epsilon$. Thus $(X,d)$ satisfies the anti-$\mathfrak{R}$-transmissible property. We next prove that $\mathfrak{R}$ is singular. For each $(n,m)\in\mathbb{N}^{2}$ and for each $\epsilon\in(0,\infty)$, we put
$(R,d_{R})=\left(L_{n},\frac{\epsilon}{\delta_{d_{n}}(L_{n})}\cdot d_{n}\right).$
Then we have $\delta_{d_{R}}(R)=\epsilon$, and
$\phi_{R}^{(n,m),R}\left(\{f_{n,l}\}_{l=1}^{\operatorname{card}(L_{n})},\delta_{d_{n}}(L_{n})/\epsilon,d_{R}\right)=0\not\in F_{R}(n,m).$
Therefore $\mathfrak{R}$ is singular. This completes the proof. ∎
Since every compact metric space is arbitrarily approximated by members of $\mathscr{F}$ in the sense of Gromov–Hausdorff, we obtain:
Proposition 4.14.
A metric space $(X,d)$ has rich pseudo-cones if and only if $\operatorname{\mathscr{P}}(X,d)$ contains all compact metric spaces.
From Theorem 1.2, we deduce the following:
Theorem 4.15.
For every non-discrete metrizable space $X$, the set of all metrics $d\in\mathrm{Met}(X)$ for which $(X,d)$ has rich pseudo-cones is dense $G_{\delta}$ in $\mathrm{Met}(X)$.
Remark 4.4.
Chen and Rossi [9] introduced the notion of locally rich compact metric spaces. They investigated the distribution of locally rich metric spaces in a space of compact metric spaces with respect to the Gromov–Hausdorff distance, and they also studied this subject in the Euclidean setting in a space of compact subspaces.
4.4. Metric inequality
Let $f\colon\mathbb{R}^{\binom{n}{2}}\to\mathbb{R}$ be a continuous function. We say that a metric space $(X,d)$ satisfies the $(n,f)$-metric inequality if for all $n$ points $a_{1},\dots,a_{n}$ in $X$ we have $f(\{d(a_{i},a_{j})\}_{i\neq j})\geq 0$. We say that a function $f\colon\mathbb{R}^{\binom{n}{2}}\to\mathbb{R}$ is positively sub-homogeneous if there exists $s\in(0,\infty)$ such that for every $x\in\mathbb{R}^{\binom{n}{2}}$ and for every $c\in(0,\infty)$ we have $f(r\cdot x)\leq r^{c}f(x)$.
Proposition 4.16.
For $n\in\mathbb{N}$, let $f\colon\mathbb{R}^{\binom{n}{2}}\to\mathbb{R}$ be a continuous function. Then satisfying the $(n,f)$-metric inequality on metric spaces is a transmissible property. Moreover, if $f$ is positively sub-homogeneous, and if there exists a metric space that does not satisfy the $(n,f)$-metric inequality, then satisfying the $(n,f)$-metric inequality on metric spaces is a transmissible property with a singular transmissible parameter.
Proof.
Let $Q=\{1\}$ and define a map $F\colon Q\to\mathcal{F}(\mathbb{R})$ by $F(1)=[0,\infty)$. Define a map $G\colon Q\to\mathcal{P}^{*}(\mathbb{N})$ by $G(1)=\{n\}$. For each metrizable space $X$, we define a map $\phi^{1,X}\colon\operatorname{Seq}(\{n\},X)\times\{1\}\times\mathrm{Met}(X)\to\mathbb{R}$ by
$\phi^{1,X}(\{a_{i}\}_{i=1}^{n},1,d)=f(\{d(a_{i},a_{j})\}_{i\neq j}).$
Let $\mathfrak{G}=(\{1\},\mathbb{R},F,G,\{1\},\phi)$. Then $\mathfrak{G}$ is a transmissible parameter.
We next show the latter part. Since there exists a metric space that does not satisfy the $(n,f)$-metric inequality, there exists a metric space $(S,d_{S})$ with $\operatorname{card}(S)=n$ that does not satisfy the $(n,f)$-metric inequality. Let $c\in(0,\infty)$ be a positive number such that for every $x\in\mathbb{R}^{\binom{n}{2}}$, and for every $r\in(0,\infty)$ we have $f(r\cdot x)\leq r^{c}f(x)$. Let $S=\{s_{i}\}_{i=1}^{n}$ and assume that $f(\{d_{S}(s_{i},s_{j})\}_{i\neq j})<0$. For every $\epsilon\in(0,\infty)$, put $(R,d_{R})=(S,\epsilon\cdot d_{S})$. Thus we have $\delta_{d_{R}}(R)=\epsilon$, and
$\phi^{1,R}(\{s_{i}\}_{i=1}^{n},1,d_{R})=f(\{\epsilon\cdot d_{S}(s_{i},s_{j})\})=\epsilon^{c}f(\{d_{S}(s_{i},s_{j})\})<0.$
This implies that $\mathfrak{G}$ is singular. This finishes the proof. ∎
Combining Theorem 1.2, Corollary 4.4 and Proposition 4.16, we obtain the following corollary:
Corollary 4.17.
Let $X$ be a non-discrete metrizable space. For $n\in\mathbb{N}$, let $f\colon\mathbb{R}^{\binom{n}{2}}\to\mathbb{R}$ be a continuous function. If $f$ is positively sub-homogeneous, and if there exists a metric space not satisfying the $(n,f)$-metric inequality, then the set of all metrics $d$ in $\mathrm{Met}(X)$ for which the space $(X,d)$ does not satisfy the $(n,f)$-metric inequality is dense open in $\mathrm{Met}(X)$.
We say that a metric space $(X,d)$ satisfies the ultratriangle inequality if for all three points $a_{1},a_{2},a_{3}$ in $X$ we have
$d(a_{1},a_{3})\leq\max\{d(a_{1},a_{2}),d(a_{2},a_{3})\}.$
Proposition 4.18.
Define a function $f\colon\mathbb{R}^{\binom{3}{2}}\to\mathbb{R}$ by
$f(x)=\max\{x_{1,2},x_{2,3}\}-x_{1,3}.$
Then the ultrametric inequality on metric spaces is equivalent to the $(3,f)$-metric inequality, and $f$ is positively sub-homogeneous.
We say that a metric space $(X,d)$ satisfies the Ptolemy inequality if for all four points $a_{1},a_{2},a_{3},a_{4}$ in $X$ we have
$d(a_{1},a_{3})d(a_{2},a_{4})\leq d(a_{1},a_{2})d(a_{3},a_{4})+d(a_{1},a_{4})d(a_{2},a_{3}).$
Proposition 4.19.
Define a function $f\colon\mathbb{R}^{\binom{4}{2}}\to\mathbb{R}$ by
$f(x)=x_{2,3}x_{1,4}+x_{1,2}x_{3,4}-x_{1,3}x_{2,4}.$
Then the Ptolemy inequality on metric spaces is equivalent to the $(4,f)$-metric inequality, and $f$ is positively sub-homogeneous.
Gromov [13] introduced the cycle condition for metric spaces as follows: Let $m\in\mathbb{N}$ and $\kappa\in\mathbb{R}$. Let $(M(\kappa),d_{M(\kappa)})$ be the two-dimensional space form of constant curvature $\kappa$. We say that a metric space $(X,d)$ satisfies the $\operatorname{Cycl}_{m}(\kappa)$ condition if for every map $f\colon\mathbb{Z}/m\mathbb{Z}\to X$ there exists a map $g\colon\mathbb{Z}/m\mathbb{Z}\to M(\kappa)$ such that
  1. (1)
    for all $i\in\mathbb{Z}/m\mathbb{Z}$, we have
    $d_{M(\kappa)}(g(i),g(i+1))\leq d(f(i),f(i+1));$
  2. (2)
    for all $i,j\in\mathbb{Z}/m\mathbb{Z}$ with $i-j\neq\pm 1$, we have
    $d_{M(\kappa)}(g(i),g(j))\geq d(f(i),f(j)),$
    where the symbol $^{\prime\prime}$ stands for the addition of $\mathbb{Z}/m\mathbb{Z}$.
Proposition 4.20.
For every $m\in\mathbb{N}$, the $\operatorname{Cycl}_{m}(0)$ condition can be represented by an $(m,C)$-metric inequality for some positively sub-homogeneous function $C$.
Proof.
For a map $g\colon\mathbb{Z}/m\mathbb{Z}\to\mathbb{R}^{2}$, we define two functions $C_{1,g},C_{2,g}:\mathbb{R}^{\binom{m}{2}}\to\mathbb{R}$ by
$C_{1,g}(x)=\min_{i\in\mathbb{Z}/m\mathbb{Z}}\{x_{i,i+1}-d_{M(0)}(g(i),g(i+1))\},$
$C_{2,g}(x)=\min_{i,j\in\mathbb{Z}/m\mathbb{Z},\ i-j\neq\pm 1}\{d_{M(0)}(g(i),g(j))-x_{i,j}\}.$
We define a function $C\colon\mathbb{R}^{\binom{m}{2}}\to\mathbb{R}$ by
$C(x)=\sup_{g:\mathbb{Z}/m\mathbb{Z}\to M(0)}\{C_{1,g}(x),\ C_{2,g}(x)\}.$
Then $C$ is continuous. For every $r\in(0,\infty)$ we have
$C(r\cdot x)$ $=\sup_{g:\mathbb{Z}/m\mathbb{Z}\to M(0)}\{C_{1,g}(r\cdot x),\ C_{2,g}(r\cdot x)\}$
$=r\cdot\sup_{g:\mathbb{Z}/m\mathbb{Z}\to M(0)}\{C_{1,g/r}(r\cdot x),\ C_{2,g/r}(r\cdot x)\}$
$=r\cdot\sup_{g:\mathbb{Z}/m\mathbb{Z}\to M(0)}\{C_{1,g}(x),\ C_{2g}(x)\}.$
Thus the function $C$ is positively sub-homogeneous.
If $m$-many points $a_{1},\dots,a_{m}$ in $X$ satisfy $C(\{d(a_{i},a_{j})\}_{i\neq j})\geq 0$, then there exists a map $g\colon\mathbb{Z}/m\mathbb{Z}\to M(0)$ such that $C_{1,g}(\{d(a_{i},a_{j})\}_{i\neq j})\geq 0$ and $C_{2,g}(\{d(a_{i},a_{j})\}_{i\neq j})\geq 0$. These two inequalities are equivalent to the conditions (1) and (2) in the $\operatorname{Cycl}_{m}(0)$ condition, respectively. Therefore the $\operatorname{Cycl}_{m}(0)$ condition is equivalent to the $(m,C)$-metric inequality. ∎
4.5. Gromov hyperbolicity
Gromov [12] introduced the notion of the Gromov hyperbolicity. We say that $(X,d)$ is Gromov hyperbolic if there exists $\delta\in[0,\infty)$ such that for all four points $a_{1},a_{2},a_{3},a_{4}$ in $X$ we have
$d(a_{1},a_{3})+d(a_{2},a_{4})\\ \leq\max\{d(a_{1},a_{2})+d(a_{3},a_{4}),d(a_{1},a_{4})+d(a_{2},a_{3})\}+2\delta,$
Proposition 4.21.
The Gromov hyperbolicity on metric spaces is equivalent to a transmissible parameter.
Proof.
Let $Q=\mathbb{Q}_{\geq 0}$. For each $\delta\in Q$, define a function $g_{\delta}\colon\mathbb{R}^{\binom{4}{2}}\to\mathbb{R}$ by
$g_{\delta}(x)=\max\{x_{1,2}+x_{3,4},x_{1,4}+x_{2,3}\}+2\delta-(x_{1,3}+x_{2,4}).$
We also define a map $F\colon Q\to\mathcal{F}(\mathbb{R})$ by $F(\delta)=[0,\infty)$. Define a map $G\colon Q\to\mathcal{P}^{*}(\mathbb{N})$ by $G(\delta)=\{4\}$. For each metrizable space $X$, we define a map $\phi^{\delta,X}\colon\operatorname{Seq}(\{4\},X)\times\{1\}\times\mathrm{Met}(X)\to\mathbb{R}$ by
$\phi^{\delta,X}(\{a_{i}\}_{i=1}^{n},1,d)=g_{\delta}(\{d(a_{i},a_{j})\}_{i\neq j}).$
Let $\mathfrak{G}=(Q,\mathbb{R},F,G,\{1\},\phi)$. Then $\mathfrak{G}$ is a transmissible parameter. Under this notations, we observe that $(X,d)$ is Gromov hyperbolic if and only if it satisfies the $\mathfrak{G}$-transmissible property. ∎
Since for every metrizable space $X$ the set of all bounded metrics in $\mathrm{Met}(X)$ is open in $\mathrm{Met}(X)$, and since every bounded metric space is Gromov hyperbolic, we obtain the following:
Proposition 4.22.
The Gromov hyperbolicity on metric spaces is not equivalent to any transmissible property with a singular transmissible parameter.
5. Local Transmissible properties
In this section, we prove Theorem 1.3. The following lemma plays a key role in the proof of Theorem 1.3.
Lemma 5.1.
For every second-countable locally compact metrizable space $X$, the space $\mathrm{Met}(X)$ is completely metrizable. In particular, it is Baire.
Proof.
Let $C(X^{2})$ be the set of all real-valued continuous functions on $X^{2}$. We define a metric $\mathcal{E}$ on $C(X^{2})$ by
$\mathcal{E}(f,g)=\min\left\{1,\sup_{(x,y)\in X^{2}}|f(x,y)-g(x,y)|\right\}.$
Note that the metric $\mathcal{E}|_{\mathrm{Met}(X)^{2}}$ on $\mathrm{Met}(X)$ generates the same topology as $\mathcal{D}_{X}$ on $\mathrm{Met}(X)$, which coincides with the uniform topology.
Note that the space $(C(X^{2}),\mathcal{E})$ is completely metrizable. By Lemma 2.8, to prove the lemma, it suffices to show that $\mathrm{Met}(X)$ is $G_{\delta}$ in $(C(X^{2}),\mathcal{E})$. We denote by $\mathrm{PM}$ the set of all $f\in C(X^{2})$ such that
  1. (1)
    for every $x\in X$ we have $f(x,x)=0$;
  2. (2)
    for all $x,y\in X$ we have $f(x,y)=f(y,x)$ and $f(x,y)\geq 0$;
  3. (3)
    for every triple $x,y,z\in X$ we have $f(x,y)\leq f(x,z)+f(z,y)$.
Namely, $\mathrm{PM}$ is the set of all continuous pseudo-metrics on $X$. Note that $\mathrm{PM}$ is a closed subset in the metric space $(C(X^{2}),\mathcal{E})$. Since all closed subsets of a metric space are $G_{\delta}$ in the whole space, the set $\mathrm{PM}$ is $G_{\delta}$ in $(C(X^{2}),\mathcal{E})$.
Using the assumption that $X$ is second-countable and locally compact, now we take a sequence $\{B(n)\}_{n\in\mathbb{N}}$ of compact subsets of $X^{2}$ with $\bigcup_{n\in\mathbb{N}}B(n)=X^{2}\setminus\Delta_{X}$, where $\Delta_{X}$ is the diagonal set of $X^{2}$, and take a sequence $\{K(n)\}_{n\in\mathbb{N}}$ of compact subsets of $X$ with $K(n)\subset\operatorname{INT}(K(n+1))$ and $\bigcup_{n\in\mathbb{N}}K(n)=X$, where $\operatorname{INT}$” means the interior.
For every $n\in\mathbb{N}$, let $\mathrm{L}_{n}$ be the set of all $f\in C(X^{2})$ for which there exists $c\in(0,\infty)$ such that for each $s>n$ we have
$\inf\{\,f(x,y)\mid x\in K(n),\ y\in X\setminus K(s)\,\}>c.$
For each $n\in\mathbb{N}$, let $\mathrm{E}_{n}$ be the set of all $f\in C(X^{2})$ such that for each $(x,y)\in B(n)$ we have $0<f(x,y)$. Note that each $\mathrm{L}_{n}$ and each $\mathrm{E}_{n}$ are open subsets of $(C(X^{2}),\mathcal{E})$.
We next prove that
$\mathrm{Met}(X)=\mathrm{PM}\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}\right)\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}\right).$
To prove $\mathrm{Met}(X)\subset\mathrm{PM}\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}\right)\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}\right)$, take $d\in\mathrm{Met}(X)$. Since $d$ is a metric on $X$, we have $d\in\mathrm{PM}$ and $d\in\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}$. To show $d\in\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}$, take an arbitrary number $n\in\mathbb{N}$. Since $K(n)\subset\operatorname{INT}(K(n+1))$ and since $d\in\mathrm{Met}(X)$, for every $s>n$ we have
$0$ $<d(K(n),X\setminus\operatorname{INT}(K(n+1)))\leq d(K(n),X\setminus K(n+1))$
$\leq d(K(n),X\setminus K(s)).$
Thus $d\in\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}$. Hence we obtain
$\mathrm{Met}(X)\subset\mathrm{PM}\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}\right)\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}\right).$
Next we take $d\in\mathrm{PM}\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}\right)\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}\right)$. Since $d\in\mathrm{PM}\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}\right)$, the function $d$ is continuous on $X^{2}$ and it is a metric on $X$. Fix $e\in\mathrm{Met}(X)$. We show that the metric $d$ is topologically equivalent to $e$. Since $d$ is continuous on $X^{2}$, the metric $d$ generates a weaker topology than that of $(X,e)$. Namely, if $\lim_{n\to\infty}x_{n}=a$ in $(X,e)$, then $\lim_{n\to\infty}x_{n}=a$ in $(X,d)$. Assume next that $\lim_{n\to\infty}x_{n}=a$ in $(X,d)$. Our purpose is to prove that $\lim_{n\to\infty}x_{n}\to a$ in $(X,e)$. Since $\{K(n)\}_{n\in\mathbb{N}}$ is a covering of $X$, there exists $s\in\mathbb{N}$ such that $a\in K(s)$. For the sake of contradiction, suppose that there exist infinitely many $i\in\mathbb{N}$ with $(X\setminus K(i))\cap\{\,x_{n}\mid{n\in\mathbb{N}}\,\}\neq\emptyset$. Then we have
$\liminf_{i\to\infty}d(K(s),X\setminus K(i))\leq\lim_{j\to\infty}d(a,x_{j})=0.$
This contradicts the fact that $d\in\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}$. Hence there exists a sufficiently large number $m\in\mathbb{N}$ such that the sequence $\{x_{n}\}_{n\in\mathbb{N}}$ is contained in $K(m)$. We may assume that $s\leq m$, which implies that $a\in K(m)$. By the compactness of $K(m)$ and by the fact that $(X,d)$ is Hausdorff, the map $1_{K(m)}\colon(K(m),e)\to(K(m),d)$ becomes a homeomorphism. Thus the restricted metrics $e|_{K(m)^{2}}$ and $d|_{K(m)^{2}}$ generate the same topology on $K(m)$. Since $\lim_{n\to\infty}x_{n}=a$ in $(X,d)$ and since $\{x_{n}\}_{n\in\mathbb{N}}$ is contained in $K(m)$, we see that $\lim_{n\to\infty}x_{n}=a$ in $(X,e)$.
As a result, the metric $d$ generates the same topology as $e$, and hence
$\mathrm{Met}(X)\supset\mathrm{PM}\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{L}_{n}\right)\cap\left(\bigcap_{n\in\mathbb{N}}\mathrm{E}_{n}\right).$
Therefore we conclude that $\mathrm{Met}(X)$ is a $G_{\delta}$ subset of $(C(X^{2}),\mathcal{E})$. ∎
Proof of Theorem 1.3.
Let $X$ be a second-countable, locally compact locally non-discrete space, and let $\mathfrak{G}$ be a singular transmissible parameter. Let $\mathfrak{G}=(Q,P,F,G,Z,\phi)$, and let $S$ be the set of all metrics $d\in\mathrm{Met}(X)$ for which $(X,d)$ satisfies the local anti-$\mathfrak{G}$-transmissible property. Take a countable open base $\{U_{i}\}_{i\in\mathbb{N}}$ of $X$, and take a family $\{R_{i}\}_{i\in\mathbb{N}}$ of $(\omega_{0}+1)$-subspaces of $X$ with $R_{i}\subset U_{i}$. Since $\{U_{i}\}_{i\in\mathbb{N}}$ is an open base of $X$, according to Lemma 4.1, we have
$S=\bigcap_{i\in\mathbb{N}}\bigcap_{q\in Q}\bigcup_{z\in Z}\bigcup_{a\in\operatorname{Seq}(G(q),U_{i})}S(X,\mathfrak{G},q,a,z).$
Corollary 4.4 implies that $S$ is $G_{\delta}$ in $\mathrm{Met}(X)$. For every $i\in\mathbb{N}$, the set
$\bigcap_{q\in Q}\bigcup_{z\in Z}\bigcup_{a\in\operatorname{Seq}(G(q),U_{i})}S(X,\mathfrak{G},q,a,z)$
contains $T(X,R_{i},\mathfrak{G})$. Proposition 4.6 implies that each $T(X,R_{i},\mathfrak{G})$ is dense in $\mathrm{Met}(X)$. By Lemma 5.1, the space $\mathrm{Met}(X)$ is Baire, and hence $S$ is dense $G_{\delta}$ in $\mathrm{Met}(X)$. This completes the proof. ∎